solve for x: 3(x-4)-1=2x+5
Answer by CharStar(110)

Are you asking for this problem to be graphed also?
3(x-4)-1=2x+5
3x-12-1=2x+5
3x-13=2x+5
Add 13 to the right side
Subtract 2x from the left side
3x-2x=5+13
x=18
Hope that helps you